• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sulkosky

    Log entry time 11:51:50 on June26,2003

    Entry number 103506

    keyword=Target Status

    I noticed that the NMR signal was less than 2 mV after 9 hours of pumping. I checked the lasers to make sure they were on. RTD 6 and 7 are colder than I expected, and the spectrum from the spectrum analyzer was far from 795 nm. I found that Lasers 6 and 7 were not at full power. So I changed the set current to 38.1 A and 40.0 A respectively. RTD 6 and 7 now are at ~ 192 deg. C and 177 deg. C respectively, which is where they should be. I expect the spin-up of Priapus to look better now.